OpenStudy (anonymous):

Can someone please help me with BASES! I need to find a base that makes this statement true: 135b+444b=601b. Where do I even start?! :(

OpenStudy (anonymous):

would making them all base 10 work?

OpenStudy (anonymous):

b represent the base

OpenStudy (mrhoola):

perhaps write the expression as log base 135 (x) + log base 444 (x) = log base 601 (x) . I wrote x so that they are all related , then try changing the base to 10 , from there try some log manipulation . I would not know how else to appraoch this problem

OpenStudy (anonymous):

i dont understand the process though, like once i do what you suggested, how do i convert them to base 10?

OpenStudy (mrhoola):

log base 135 (x) = log base 10 (x) / log base 10 (135) That is what I found on google . but you might want to double check. this approach may be the incorrect one . nonethe less its a thought

OpenStudy (anonymous):

im not familiar with bases at all but ill try! thank you!

OpenStudy (anonymous):

@ganeshie8 can you please help me with this problem?

OpenStudy (mrhoola):

applying the change in base that would mean you would also have to find a commonality that between the 136 , 444 , and 601 . In base 10 I dont think those values are acceptable , you might have try try a different base

ganeshie8 (ganeshie8):

135 444 ----- 601 look at the units place addition 5+4 = 11

ganeshie8 (ganeshie8):

\[(5+4)_b = (11)_b\] \[ 9 = 1*b^1 + 1*b^0\] \[ 9 = b+1\] \[ 8 = b\]

OpenStudy (anonymous):

so does that mean that base 8 makes the entire statement true?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

how did you know to do that?

ganeshie8 (ganeshie8):

13 `5` 44 `4` ----- 60 `1`

ganeshie8 (ganeshie8):

Notice that adding the last digits is producing a carry : `1` 13 `5` 44 `4` ----- 60 `1`

ganeshie8 (ganeshie8):

that means \((5+4)_b = (11)_b\) did u get this far ?

OpenStudy (anonymous):

so thats just (11)b=11b?

OpenStudy (anonymous):

oh no! (9)b=(11)b?

ganeshie8 (ganeshie8):

in the next step we're changing everything to base 10

ganeshie8 (ganeshie8):

5+4 = 9 in base 10

ganeshie8 (ganeshie8):

\((5+4)_b = (11)_b\) \((9)_{10} = 1\times b^1 + 1\times b^0\)

ganeshie8 (ganeshie8):

you must be knowing converting a different base number to base10 ?

OpenStudy (anonymous):

i don't have much experience with bases :( thats what I'm trying to figure out

ganeshie8 (ganeshie8):

Lets start from scratch

ganeshie8 (ganeshie8):

so you're comfortable only with base10 ?

OpenStudy (anonymous):

yes! i am barely learning base 10

ganeshie8 (ganeshie8):

the main reason humans use 10 symbols to count is because they have 10 fingers

OpenStudy (anonymous):

makes sense! haha

ganeshie8 (ganeshie8):

An alien with only 5 fingers might feel comfortable using base5 instead : \( (4 + 4)_{5} = (13)_5\)

ganeshie8 (ganeshie8):

lets see how the alien is getting 13 for that addition in base5

ganeshie8 (ganeshie8):

suppose we give the alien \(8\) stones and ask it to count : 

OpenStudy (anonymous):

so 1 set and 3 extra

ganeshie8 (ganeshie8):

it starts counting up to 5 using its 5 fingers and puts it aside and says it is one `five`: